# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ 5.1 Derivatives of the Matérn covariance kernel
The differentiability of Matérn covariance functions are well-known [55, 22]. However, the Lipschitz continuity result, given in equation (25) below, is somewhat more obscure, with only the $\alpha\not\in\mathbb{N}$ case considered in [51], without proofs. For the sake of clarity and completeness, we collect all these results here and provide the corresponding proofs in the supplement.
@@
Lemma 4 shows that the Matérn covariance function $K(\cdot,\cdot)$, viewed as a map from $\mathbb{R}^{2d}\to\mathbb{R}$, is $2\underline{\alpha}$ times differentiable, with Lipschitz continuous $2\underline{\alpha}$th order derivatives. The Matérn covariance function belongs to $C^{\alpha}(\mathbb{R}^{d})\times C^{\alpha}(\mathbb{R}^{d})$ in the sense that for integer $\alpha$ the function is weakly differentiable of order $(k,k),\forall|k|=\alpha$. However, its Hölder-$(\alpha,\alpha)$ norm is infinite since the weak derivatives of order $(k,k),\forall|k|=\alpha$ are not uniformly bounded. Moreover, the sample paths of the process are almost surely $\underline{\alpha}$ times differentiable with their $\underline{\alpha}$th derivatives being almost surely Lipschitz continuous of order $\alpha-\underline{\alpha}$, up to a logarithmic term.
The smoothness properties described in Lemma 4 carry over to the covariance kernels of the conditional processes as well. More precisely, for an arbitrary finite set $A=\{w_{1},w_{2},\cdots,w_{m}\}\subset\mathbb{R}^{d}$, let us define the process $\tilde{Z}$ as
@@
The process $\tilde{Z}$ can be regarded as the Gaussian process conditioned on the set $A$. We show below that it inherits several key properties of the mother Matérn process $Z$. Remarkably, these properties do not depend on the conditioning set $A$.
@@
All constant multipliers in the upper bounds are independent of $x,h\in\mathbb{R}^{d}$ and the set $A$.
The universality of the constants in Lemma 5 across the different conditioning sets $A$ is implied by the fact that the variance of the conditional process $\tilde{Z}$ does not exceed the variance of its mother GP $Z$. In view of the relationship between the derivatives of GPs and the derivatives of the corresponding covariance functions established in Proposition I.3 of [22], the above intuition applies also to the derivatives of the GPs $Z$ and $\tilde{Z}$, up to order $\underline{\alpha}$. Therefore, regardless of the conditioning set $A$, the conditional process $\tilde{Z}$ inherits similar smoothness properties as the mother GP $Z$. This feature plays a crucial role in the analysis of the conditional distribution in the next sections. We also note that the logarithmic term in the previous lemmas for $\alpha\in\mathbb{N}$ follows from the singularity of the modified Bessel functions of second kind at integer parameters. In the rest of the paper, for simplicity of notations, we use $\ln(1/\|h\|_{2})]^{\mathbbm{1}_{\{\alpha\in\mathbb{N}\}}}$ to denote that the logarithmic term only occurs for integer $\alpha$ values.
@@ 5.2 Conditional distribution
The conditional distribution formulas (8) and (2.2) defining the Vecchia GPs are the starting point of our theoretical analysis. In this section, we study three aspects of these conditional distributions: the variance in equation (2.2), the conditional expectation in equation (2.2), and the recursive application of this conditional expectation formula.
@@ 5.2.1 Variance and conditional expectation
Under the Norming condition 2, the distance among vertices in the set $\mathrm{pa}(X_{i})\cup\{X_{i}\}$ is bounded by $\gamma^{-\eta(i)}$ up to a constant multiplier, where $\eta(i)=j$ denotes the layer allocation of $X_{i}\in\mathcal{N}_{j}$. We recall, that the observations are ordered based on their layer allocations, hence the functions $i\rightarrow\eta(i)$ is monotone increasing. We investigate the asymptotic behavior of the Gaussian random variable $Z^{\tau,s}_{X_{i}}-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]$ as $i$ goes to infinity. The following theorem provides lower and upper bounds for its variance, which by equations (2.2) and (2.2) is the same as the variance of the conditional Matérn GP process.
@@
Let the mother Gaussian process $Z^{\tau,s}$ be the rescaled Matérn process with smoothness parameter $\alpha$. Suppose Condition 2 is satisfied, then for all $X_{i}\in\mathcal{N}_{j}$, j=1,,Jj=1,...,J, we have
@@
Theorem 1 plays an essential role in determining the small ball probabilities of Vecchia GPs. Specifically, for a finite training set $\mathcal{X}_{n}$, the number of layers $\eta(n)$ satisfies $\gamma^{d\eta(n)}\asymp n$. Therefore, for $X_{i}\in\mathcal{N}_{\eta(n)}$ and $\tau=s=1$, $\mathrm{Var}\big{\{}Z_{X_{i}}-\mathbb{E}[{Z}_{X_{i}}|{Z}_{\mathrm{pa}(X_{i})}]\big{\}}\asymp n^{-2\alpha/d}.$ In other words, the variance at $X_{i}\in\mathcal{N}_{\eta(n)}$ goes to zero with a polynomial rate in $n$, where the order is determined by the smoothness parameter $\alpha$ of the mother Matérn process. Hence smoother mother GPs posses faster decaying variances (29).
We proceed to study the conditional expectation $\mathbb{E}[\hat{Z}^{\tau,s}_{X_{i}}|\hat{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]$ in equation (2.2), which is the Gaussian interpolation, also called kriging in spatial statistics literature [20], of $\hat{Z}^{\tau,s}_{X_{i}}$ using the random vector $\hat{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}$. Recalling Condition 2, for all $i\geq i_{0}$ and $X_{i}\in\mathcal{N}_{\eta(i)}$, its parent set $\mathrm{pa}(X_{i})$ is contained in a cube of side length $c_{L}\gamma^{-\eta(i)}$ and $\mathrm{pa}(X_{i})$ is a norming set in this cube with norming constant $c_{N}$. Therefore, given arbitrary $X_{i}$, there exists a norming set $A=\{w_{i},1\leq i\leq m\}\subset[0,1]^{d}$ with norming constant $c_{N}$ and $x^{\prime}\in\mathbb{R}^{d}$, such that for all rescaling parameter $\tau>0$,
@@
Then, in view of the stationarity of the Matérn process, the conditional expectation in (2.2) can be written as
@@
$=K_{\nu A,\nu x^{*}}^{T}K_{\nu A,\nu A}^{-1}Z^{\tau,s}_{\mathrm{pa}(X_{i})},$ (30)
where $\nu=\tau c_{L}\gamma^{-\eta(i)}\asymp\tau\gamma^{-\eta(i)}$. Note that the conditional expectation above depends on the covariance function $K$, characterizing the spatial dependencies of the process $Z$. However, we show below, that under certain regularity conditions, for fixed $x^{*}$ and $A$, if letting $\nu$ tend to zero, the limit for the Gaussian interpolation weights $K_{\nu A,\nu A}^{-1}K_{\nu A,\nu x^{*}}$ is free of the covariance function. This limit is often referred to as the flat limit in the literature of radial basis functions [15, 46, 35]. For Matérn covariance functions, the following lemma shows that the flat limit coincides with the polynomial interpolation weights $V_{A}^{-1}v_{x}^{*}$ if the set $A$ is a norming set with respect to an appropriately chosen families of polynomials.
@@
Let $A$ be a norming set on $\mathscr{P}_{\underline{\alpha}}([0,1]^{d})$ with norming constant $c_{N}$. Suppose $x\in[0,1]^{d}$, then for all $\nu<1$,
@@
where $V_{A}$ and $v_{x}$ are defined in (14) and (15). Similarly, for $\nu,\nu_{0}\in(0,+\infty)$ and $|\nu-\nu_{0}|\leq 1$,
@@
Lemma 6 provides an example of flat limits of radial basis function interpolation, a topic with extended literature, studied over twenty years, see for instance [15, 35]. The general idea is that for smooth enough kernels (i.e. Matérn covariance functions in our case) and unisolvent set $A$ with respect to some family of polynomials, the limit of equation (30) is equal to the limit of polynomial interpolation within that family. With this in mind, we discuss the conditions of Lemma 6. For $A$ to be a norming set of a polynomial space, its cardinality has to be at least the dimension of this polynomial space. The covariance function $K(\cdot,\cdot)$ of the Matérn process with regularity $\alpha$ belongs to $C^{2\underline{\alpha}}$, i.e., it is $\underline{\alpha}$ times differentiable in both covariates. Intuitively, the polynomial space describing the local behavior of Matérn covariance function is of order $\underline{\alpha}$, which requires the set $A$ to be a norming set on $\mathscr{P}_{\underline{\alpha}}([0,1]^{d})$.
In view of Lemma 6, the norming constant $c_{N}$ appears in the upper bounds, hence it is important to have a tight control on it. However, as discussed next, this is a highly challenging task in general. Recall, that $A$ is a norming set if it is a unisolvent set with respect to some polynomials. In the unidimensional case, a set $A$ with $m$ distinct elements is always unisolvent with respect to the polynomials up to (m1)(m-1)th order. In higher dimensions $d\geq 2$, however, it is possible that the elements of $A$ belong to a lower dimensional polynomial manifold and, as a consequence, the corresponding Vandermonde matrix $V_{A}$ is singular. In such cases, the limit of Gaussian interpolation on the set $A$ is interpreted within a family of lower order polynomials. We note that the assumption of $A$ being exactly a unisolvent set of $\mathscr{P}_{l}([0,1]^{d})$ is not necessary. The limit of the expression (30) still exists when choosing arbitrary ${l+d\choose l}\leq m<{l+1+d\choose l+1}$. However, the corresponding, asymptotically equivalent formulation is more complex, depending on the derivatives of the covariance function at the origin, see [35]. For instance, if $A$ is a strict superset of a norming set and $\alpha$ is an even integer then the limit is the polyharmonic spline interpolation, see [46]. We are not aware of any result addressing this limit in a more general setting. Therefore, for simplicity, we consider unisolvent sets from now on.
@@ 5.2.2 Recursive interpolation
Another important aspect of Vecchia GPs is the repeated application of the conditional expectation formula (30). Specifically, for all $X_{i}\in\mathcal{X}$, we denote the double parent set of $X_{i}$ as $\mathrm{pa}^{2}(X_{i})\triangleq\cup_{X\in\mathrm{pa}(X_{i})}\mathrm{pa}(X)$. However, despite the conditional distribution of $Z^{\tau,s}_{X_{i}}$ and its Vecchia approximation $\hat{Z}^{\tau,s}_{X_{i}}$ are the same given the processes at the corresponding parent set $\mathrm{pa}(X_{i})$, this is (typically) not the case when we condition on the processes at the double parent set $\mathrm{pa}^{2}(X_{i})$. In other words, applying the conditional expectation formula of Vecchia Gaussian processes more than once will result in a different conditional distribution than from the mother GP.
Therefore, to formally study the properties of (repeated) Gaussian interpolation we introduce the operator $G_{j}$ that maps a function $f:\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}\to\mathbb{R}$ to a function $G_{j}(f):\cup_{j^{\prime}=0}^{j+1}\mathcal{N}_{j^{\prime}}\to\mathbb{R}$, such that
@@
where $\bm{f}(\mathrm{pa}(X))$ denotes the vector $(f(x))_{x\in\mathrm{pa}(X)}$, see Section 1.3.3. The above operator provides the Gaussian interpolation at locations $X\in\mathcal{N}_{j+1}$ given the function values at the lower layers $\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}$. In view of Lemma 6, the limit of the Gaussian interpolation is the polynomial interpolation. Let us denote the corresponding polynomial interpolation operator $P_{j}$ mapping a function $f:\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}\to\mathbb{R}$ to a function $P_{j}(f):\cup_{j^{\prime}=0}^{j+1}\mathcal{N}_{j^{\prime}}\to\mathbb{R}$ as
@@
Furthermore, let us endow the space of functions on $\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}}$ with the (empirical) supremum norm $\|f\|_{\infty,j}=\sup_{X\in\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}}}f(X)$. For the operator $P_{j}$, mapping functions from the space $\ell^{\infty}(\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}})$ to $\ell^{\infty}(\cup_{j^{\prime}=1}^{j+1}\mathcal{N}_{j^{\prime}})$, we define its (empirical) $\ell^{\infty}$ operator norm as
$\|P_{j}\|\triangleq\sup_{f\in\ell^{\infty}(\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}})}\frac{\|P_{j}(f)\|_{\infty,j+1}}{\|f\|_{\infty,j}}.$
We apply the same operator norm definition to $G_{j}$ and their compositions. Then in view of Lemma 6, the difference between the operators $G_{j}$ and $P_{j}$ tends to zero in operator norm. Therefore, to give an upper bound for the operator norm of the recursive Gaussian interpolation operator $G_{j_{2}}G_{j_{2}-1}\cdots G_{j_{1}}(f)$, for some $j_{1}<j_{2},$ it is sufficient to control the corresponding recursive interpolation with the polynomial operators.
@@
$\vartheta_{n}=\sup\left\{\frac{\|P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)\|_{\infty,j_{1}+1}}{\|f\|_{\infty,j}}:f\in\ell^{\infty}(\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}}),0\leq j_{1}\leq j_{2}\leq\eta(n)\right\}.$ (34)
Due to the close connection of Gaussian and polynomial interpolations, this term plays a crucial role for deriving small deviation bounds for Vecchia GPs. Note that it depends both on the data set $\mathcal{X}_{n}$ and the DAG built on the top of it.
@@
In our analysis, we consider grid data (19). In Lemma 14, deferred to Section B.5 of the Supplementary Material, we obtain tight control for $\vartheta_{n}$ by showing that it is bounded for grid data (19) and Layered Norming sets. Although, through considerable numerical evidence, we believe that $\vartheta_{n}=O(1)$ holds more generally, we do not have a rigorous proof for this claim. The difficulty lies in achieving a tight control on the multiple ($n$ dependent) recursive application of the operators $P_{j},j\in\mathbb{N}$, which is not implied directly by the uniform control of the individual operators. Specifically, in the considered Layered Norming DAGs, the total number of layers is $\eta(n)\asymp\ln n$. Therefore, even if we have a uniform control over the operators $\|P_{j}\|\leq c,\forall j\in\mathbb{N}$ for some constant $c>1$, this still results in a sub-optimally large upper bound for the repeated interpolation
@@
Therefore, we need to consider the composition of these operators as a whole, which is exceedingly difficult to handle without any additional structural or geometrical assumptions on $\mathcal{X}_{n}$. The case becomes considerably easier for grid data. The proof of Lemma 14 in the supplement consists of showing that in this special case the polynomial interpolation is equivalent to convolution, which can be further converted to multiplication via Fourier transform. Then by semi-explicit computations we verify that $\vartheta_{n}=O(1)$.
Then, by exploiting the close asymptotic equivalence of the polynomial and Gaussian interpolations, see Lemma 6, we bound the operator norm of the recursive Gaussian interpolation with the help of $\vartheta_{n}$ given in (34).
@@
Under Condition 2, we have for all $0\leq j_{1}\leq j_{2}$,
$\|G_{j_{2}}\cdots G_{j_{1}+1}G_{j_{1}}\|_{\infty}\lesssim\tilde{\vartheta}_{n},$
@@
Note that $\vartheta_{n}=O(1)$ implies that $\tilde{\vartheta}_{n}=O(1)$, which is the case for grid data. In the following, recursive Gaussian interpolation will appear in certain technical lemmas. For simplicity, in the rest of theoretical results, whenever recursive Gaussian interpolation is involved, we explicitly write out the $\tilde{\vartheta}_{n}$ term in the upper bound. Then the corresponding results can be applied beyond the grid data case as well, as soon the term $\vartheta_{n}$ or more generally $\tilde{\vartheta}_{n}$ can be controlled.
We conclude this section with some general remarks. Theorem 1 and Lemma 6 study the variance and conditional expectations that define the Vecchia GPs, while Lemma 7 provides an upper bound for the operator norm of the recursive application of Gaussian process interpolation. These three lemmas are the building blocks for deriving the small deviation bounds for GPs, controlling the metric entropy of the associated RKHS and finally, deriving posterior contraction rates for Vecchia GPs.

### v4

@@ 4.1 Derivatives of the Matérn covariance kernel
The differentiability properties of Matérn covariance functions are well-known [60, 22]. However, the Lipschitz continuity result given in equation (20) below, is less commonly discussed. In particular, only the case $\alpha\not\in\mathbb{N}$ is considered in [56], and no proofs are provided. For clarity and completeness, we collect all relevant results here and provide the corresponding proofs in the supplement.
@@
Lemma 1 shows that the Matérn covariance function $K(\cdot,\cdot)$, viewed as a map from $\mathbb{R}^{2d}\to\mathbb{R}$, is $2\underline{\alpha}$ times differentiable, with its $2\underline{\alpha}$th-order derivatives being Lipschitz continuous. The Matérn covariance function belongs to $C^{\alpha}(\mathbb{R}^{d})\times C^{\alpha}(\mathbb{R}^{d})$ in the sense that, for integer $\alpha$, it is weakly differentiable of order $(k,k),\forall|k|=\alpha$. However, its Hölder-$(\alpha,\alpha)$ norm is infinite, since the weak derivatives of order $(k,k),\forall|k|=\alpha$ are not uniformly bounded. Moreover, the sample paths of the process are almost surely $\underline{\alpha}$ times differentiable with their $\underline{\alpha}$th derivatives being almost surely Lipschitz continuous of order $\alpha-\underline{\alpha}$, up to a logarithmic term.
The smoothness properties described in Lemma 1 also carry over to the covariance kernels of the conditional processes. More precisely, for an arbitrary finite set $A=\{w_{1},w_{2},\cdots,w_{m}\}\subset\mathbb{R}^{d}$, we define the process $\tilde{Z}$ as
@@
The process $\tilde{Z}$ can be regarded as the Gaussian process conditioned on the set $A$. We show below that it inherits several key properties of the mother Matérn process $Z$. Remarkably, these properties do not depend on the choice of the conditioning set $A$.
@@
All constant multipliers in the upper bounds are independent of $x,h\in\mathbb{R}^{d}$ and the choice of the set $A$.
The universality of the constants in Lemma 2 across different conditioning sets $A$ is a consequence of the fact that the variance of the conditional process $\tilde{Z}$ does not exceed the variance of its mother GP $Z$. In view of the relationship between the derivatives of GPs and the derivatives of the corresponding covariance functions, as established in Proposition I.3 of [22], this reasoning extends to the derivatives of the GPs $Z$ and $\tilde{Z}$, up to order $\underline{\alpha}$. Therefore, regardless of the conditioning set $A$, the conditional process $\tilde{Z}$ inherits similar smoothness properties as the mother GP $Z$. This feature plays a crucial role in analyzing the conditional distribution in the subsequent sections. We also note that the logarithmic term in the preceding lemmas for $\alpha\in\mathbb{N}$ arises from the singularity of the modified Bessel functions of second kind at integer parameters. For notational simplicity, throughout the rest of the paper we write $\ln(1/\|h\|_{2})]^{\mathbbm{1}_{\{\alpha\in\mathbb{N}\}}}$ to indicate that the logarithmic term appears only when $\alpha$ is an integer.
@@ 4.2 Conditional distribution
The conditional distribution formulas (8) and (2.2) defining the Vecchia GPs form the starting point of our theoretical analysis. In this section, we study three aspects of these conditional distributions: the variance in equation (2.2), the conditional expectation in equation (2.2), and the recursive application of this conditional expectation formula.
@@ 4.2.1 Variance and conditional expectation
Under the Norming condition 2, the distance among vertices in the set $\mathrm{pa}(X_{i})\cup\{X_{i}\}$ is bounded by $\gamma^{-\eta(i)}$ up to a constant multiplier, where $\eta(i)=j$ denotes the layer allocation of $X_{i}\in\mathcal{N}_{j}$. We recall that the observations are ordered based on their layer allocations, hence the function $i\rightarrow\eta(i)$ is monotone increasing. We investigate the asymptotic behavior of the Gaussian random variable $Z^{\tau,s}_{X_{i}}-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]$ as $i$ tends to infinity. The following theorem provides lower and upper bounds for its variance, which, by equations (2.2) and (2.2), coincides with the variance of the conditional Matérn GP process.
@@
Let the mother Gaussian process $Z^{\tau,s}$ be the rescaled Matérn process with smoothness parameter $\alpha$. Suppose that Condition 2 is satisfied. Then, for all $X_{i}\in\mathcal{N}_{j}$, j=1,,Jj=1,...,J, we have
@@
Theorem 1 plays an essential role in determining the small ball probabilities of Vecchia GPs. Specifically, for a finite training set $\mathcal{X}_{n}$, the number of layers $\eta(n)$ satisfies $\gamma^{d\eta(n)}\asymp n$. Therefore, for $X_{i}\in\mathcal{N}_{\eta(n)}$ and $\tau=s=1$, $\mathrm{Var}\big\{Z_{X_{i}}-\mathbb{E}[{Z}_{X_{i}}|{Z}_{\mathrm{pa}(X_{i})}]\big\}\asymp n^{-2\alpha/d}.$ In other words, the variance at $X_{i}\in\mathcal{N}_{\eta(n)}$ decays to zero with a polynomial rate in $n$, where the order is determined by the smoothness parameter $\alpha$ of the mother Matérn process. Hence smoother mother GPs exhibit faster decaying variances, as indicated in (24).
We proceed to study the conditional expectation $\mathbb{E}[\hat{Z}^{\tau,s}_{X_{i}}|\hat{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]$ in equation (2.2), which corresponds to Gaussian interpolation, also called kriging in the spatial statistics literature [20], of $\hat{Z}^{\tau,s}_{X_{i}}$ using the random vector $\hat{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}$. Recalling Condition 2, for all $i\geq i_{0}$ and $X_{i}\in\mathcal{N}_{\eta(i)}$, its parent set $\mathrm{pa}(X_{i})$ is contained in a cube of side length $c_{L}\gamma^{-\eta(i)}$ and $\mathrm{pa}(X_{i})$ forms a norming set in this cube with norming constant $c_{N}$. Therefore, given an arbitrary $X_{i}$, there exists a norming set $A=\{w_{i},1\leq i\leq m\}\subset[0,1]^{d}$ with norming constant $c_{N}$ and a vector $x^{\prime}\in\mathbb{R}^{d}$ such that, for any rescaling parameter $\tau>0$,
@@
In view of the stationarity of the Matérn process, the conditional expectation in (2.2) can then be expressed as
@@
$=K_{\nu A,\nu x^{*}}^{T}K_{\nu A,\nu A}^{-1}\hat{Z}^{\tau,s}_{\mathrm{pa}(X_{i})},$ (25)
where $\nu=\tau c_{L}\gamma^{-\eta(i)}\asymp\tau\gamma^{-\eta(i)}$. Note that the conditional expectation above depends on the covariance kernel $K$, which characterizes the spatial dependencies of the process $Z$. However, we show below, that under certain regularity conditions, for fixed $x^{*}$ and $A$, as $\nu$ tends to zero, the limit for the Gaussian interpolation weights $K_{\nu A,\nu A}^{-1}K_{\nu A,\nu x^{*}}$ becomes independent of the covariance function. This limit is often referred to as the flat limit in the radial basis function literature [15, 49, 36]. For Matérn covariance functions, the following lemma shows that the flat limit coincides with the polynomial interpolation weights $V_{A}^{-1}v_{x}^{*}$, provided that the set $A$ is a norming set with respect to an appropriately chosen family of polynomials.
@@
Let $A$ be a norming set on $\mathscr{P}_{\underline{\alpha}}([0,1]^{d})$ with norming constant $c_{N}$. Suppose $x\in[0,1]^{d}$, then for all $\nu\leq 1$,
@@
where $V_{A}$ and $v_{x}$ are defined in (B.1) and (B.2). Similarly, for $\nu,\nu_{0}\in(0,+\infty)$ and $|\nu-\nu_{0}|\leq 1$,
@@
Lemma 3 provides an example of flat limits in radial basis function interpolation, a topic with a long-standing literature spanning over twenty years, see for instance [15, 36]. The general idea is that, for sufficiently smooth kernels (i.e. Matérn covariance functions in our case) and an unisolvent set $A$ with respect to a given family of polynomials, the limit of equation (25) coincides with the limit of polynomial interpolation within that family. With this in mind, we now discuss the conditions of Lemma 3. For $A$ to serve as a norming set of a polynomial space, its cardinality has to be at least equal to the dimension of the polynomial space. The covariance function $K(\cdot,\cdot)$ of the Matérn process with regularity $\alpha$ belongs to $C^{2\underline{\alpha}}$, i.e. it is $\underline{\alpha}$ times differentiable in each covariate. Intuitively, the polynomial space describing the local behavior of Matérn covariance function is of order $\underline{\alpha}$, which requires the set $A$ to be a norming set for $\mathscr{P}_{\underline{\alpha}}([0,1]^{d})$.
In view of Lemma 3, the norming constant $c_{N}$ appears in the upper bounds, making it important to have a tight control over it. However, as we discuss next, this is generally a highly challenging task. Recall that $A$ is a norming set if it is unisolvent with respect to a given family of polynomials. In the unidimensional case, a set $A$ with $m$ distinct elements is always unisolvent with respect to polynomials of degree up to (m1)(m-1)th. In higher dimensions ($d\geq 2$), however, it is possible for the elements of $A$ to lie on a lower-dimensional polynomial manifold, in which case the corresponding Vandermonde matrix $V_{A}$ is singular. In such cases, the limit of Gaussian interpolation on $A$ is interpreted within a family of lower-order polynomials. We note that the assumption that $A$ is exactly a unisolvent set of $\mathscr{P}_{l}([0,1]^{d})$ is not strictly necessary. The limit of the expression in (25) still exists for any choice of $m$ satisfying ${l+d\choose l}\leq m<{l+1+d\choose l+1}$. However, the corresponding, asymptotically equivalent formulation is more complex, as it depends on the derivatives of the covariance function at the origin, see [36]. For instance, if $A$ is a strict superset of a norming set and $\alpha$ is an even integer, the limit corresponds to polyharmonic spline interpolation [49]. We are not aware of any result addressing this limit in a more general setting. Therefore, for simplicity, we restrict our discussion to unisolvent sets from now on.
@@ 4.2.2 Recursive interpolation
Another important aspect of Vecchia GPs is the repeated application of the conditional expectation formula (25). Specifically, for all $X_{i}\in\mathcal{X}$, we define the double parent set of $X_{i}$ as $\mathrm{pa}^{2}(X_{i})\triangleq\cup_{X\in\mathrm{pa}(X_{i})}\mathrm{pa}(X)$. Although the conditional distributions of $Z^{\tau,s}_{X_{i}}$ and its Vecchia approximation $\hat{Z}^{\tau,s}_{X_{i}}$ coincide when conditioned on the processes at the corresponding parent set $\mathrm{pa}(X_{i})$, this is generally not the case when conditioning on the processes at the double parent set $\mathrm{pa}^{2}(X_{i})$. In other words, applying the conditional expectation formula of Vecchia Gaussian processes more than once typically produces a conditional distribution that differs from that of the mother GP.
To formally study the properties of repeated Gaussian interpolation, we introduce the operator $G_{j}$ that maps a function $f:\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}\to\mathbb{R}$ to a function $G_{j}(f):\cup_{j^{\prime}=0}^{j+1}\mathcal{N}_{j^{\prime}}\to\mathbb{R}$, such that
@@
where $\bm{f}(\mathrm{pa}(X))$ denotes the vector $(f(x))_{x\in\mathrm{pa}(X)}$, see Section 1.3.3. The operator $G_{j}$ performs Gaussian interpolation at locations $X\in\mathcal{N}_{j+1}$ using the function values at the lower layers $\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}$. In view of Lemma 3, the limit of the Gaussian interpolation corresponds to polynomial interpolation. Let us denote the associated polynomial interpolation operator by $P_{j}$, mapping a function $f:\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}\to\mathbb{R}$ to a function $P_{j}(f):\cup_{j^{\prime}=0}^{j+1}\mathcal{N}_{j^{\prime}}\to\mathbb{R}$ as
@@
Furthermore, let us endow the space of functions on $\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}}$ with the (empirical) supremum norm $\|f\|_{\infty,j}=\sup_{X\in\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}}}f(X)$. For the operator $P_{j}$, which maps functions from the space $L_{\infty}(\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}})$ to $L_{\infty}(\cup_{j^{\prime}=1}^{j+1}\mathcal{N}_{j^{\prime}})$, we define its (empirical) $L_{\infty}$ operator norm as
$\|P_{j}\|\triangleq\sup_{f\in L_{\infty}(\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}}),f\neq 0}\frac{\|P_{j}(f)\|_{\infty,j+1}}{\|f\|_{\infty,j}}.$
We define the supremem norm for $G_{j}$ and their compositions in the same way. Then, in view of Lemma 3, the difference between the operators $G_{j}$ and $P_{j}$ tends to zero in operator norm. Therefore, to obtain an upper bound for the operator norm of the recursive Gaussian interpolation operator $G_{j_{2}}G_{j_{2}-1}\cdots G_{j_{1}}(f)$, for some $j_{1}<j_{2},$ it is sufficient to control the corresponding recursive interpolation using the polynomial operators.
@@
$\vartheta_{n}=\sup\left\{\frac{\|P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)\|_{\infty,j_{1}+1}}{\|f\|_{\infty,j}}:f\in L_{\infty}(\cup_{j^{\prime}=1}^{j}\mathcal{N}_{j^{\prime}}),0\leq j_{1}\leq j_{2}\leq\eta(n)\right\}.$ (29)
Due to the close connection between Gaussian and polynomial interpolation, this term plays a crucial role in deriving small deviation bounds for Vecchia GPs. Note that it depends both on the data set $\mathcal{X}_{n}$ and on the DAG constructed over it.
@@
In our analysis, we consider grid data (14). In Lemma E.1, deferred to Section E.5 of the Supplementary Material, we obtain tight control of $\vartheta_{n}$ by showing that it is bounded for grid data (14) and layered norming sets. Although considerable numerical evidence suggests that $\vartheta_{n}=O(1)$ holds more generally, we do not have a rigorous proof of this claim. The difficulty lies in tightly controlling the multiple ($n$ dependent) recursive applications of the operators $P_{j},j\in\mathbb{N}$, which is not directly implied by the uniform control of the individual operators. Specifically, for the considered layered norming DAGs, the total number of layers satisfy $\eta(n)\asymp\ln n$. Therefore, even if we have uniform control over the operators $\|P_{j}\|\leq c,\forall j\in\mathbb{N}$ for some constant $c>1$, this still yields a suboptimally large upper bound for the repeated interpolation
@@
Therefore, it is necessary to consider the composition of these operators as a whole, which is exceedingly difficult without any additional structural or geometrical assumptions on $\mathcal{X}_{n}$. The problem becomes considerably simpler for grid data. In the proof of Lemma E.1 we show that, in this special case the polynomial interpolation is equivalent to convolution, which can be further converted to multiplication via the Fourier transform. Semi-explicit computations then verify that $\vartheta_{n}=O(1)$.
Then, by exploiting the close asymptotic equivalence of polynomial and Gaussian interpolations, see Lemma 3, we can bound the operator norm of recursive Gaussian interpolation using $\vartheta_{n}$ as defined in (29).
@@
Under Condition 2, for all $0\leq j_{1}\leq j_{2}$,
$\|G_{j_{2}}\cdots G_{j_{1}+1}G_{j_{1}}\|\lesssim\tilde{\vartheta}_{n},$
@@
Note that $\vartheta_{n}=O(1)$ implies $\tilde{\vartheta}_{n}=O(1)$, which holds for grid data. In the following, recursive Gaussian interpolation appears in several technical lemmas. For simplicity, in the rest of the remainder, whenever recursive Gaussian interpolation is involved, we explicitly include the $\tilde{\vartheta}_{n}$ term in the upper bound. This allows the results to be applied beyond the grid data case, as long as $\vartheta_{n}$ or, more generally $\tilde{\vartheta}_{n}$, can be controlled.
We conclude this section with some general remarks. Theorem 1 and Lemma 3 study the variance and conditional expectations that define the Vecchia GPs, while Lemma 4 provides an upper bound for the operator norm of the recursive Gaussian interpolation. Together, these three results form the building blocks for deriving small deviation bounds for GPs, controlling the metric entropy of the associated RKHS and ultimately deriving posterior contraction rates for Vecchia GPs.